### v2

# Multicopy Reinforcement Learning Agents

Abstract
This paper examines a novel type of multi-agent problem, in which an agent makes multiple identical copies of itself in order to achieve a single agent task better or more efficiently. This strategy improves performance if the environment is noisy and the task is sometimes unachievable by a single agent copy. We propose a learning algorithm for this multicopy problem which takes advantage of the structure of the value function to efficiently learn how to balance the advantages and costs of adding additional copies.
[…]
2695
1 Introduction
In most multi-agent domains, agents either cooperate or compete to achieve their goals [6, 4, 15]. This paper seeks to address a different type of multi-agent scenario, one in which agents duplicate themselves to execute multiple copies of the same policy to combat noise and risk in the problem domain. One such domain is packet forwarding in mobile wireless networks, where duplicate packets have been shown to improve reliability and reduce delay [10] as well as other domains, like robot search and rescue, where duplicate agents may improve performance.
We assume a task that, in a noiseless and no risk environment, could be completed by a single agent following an optimal policy. However, when noise increases to a sufficiently high level a single agent may fail to complete the task. In this case, an agent that can duplicate itself and make several attempts concurrently to achieve the goal may be more successful, or able to complete the task more quickly.
We will assume for now that (1) the only coordination between agents occurs when new copies are made at the start state, and that (2) after that point each agent copy operates completely independently of the others. We will also assume that (3) aside from the points at which copies are made the agents all follow identical policies. These three assumptions will give a particular form to the value function for the agent group, a form which factors easily and has many of the properties of a single agent value function, without requiring consideration of the full cross product of agent states and actions even in the exact tabular case.
2 Background: Reinforcement Learning
Reinforcement Learning (RL) assumes an agent existing in an environment which can be represented by a Markov Decision Process (consisting of states $S$, actions $A$, and transition probabilities between states) [12]. The agent seeks to maximize a reward signal over sequential, extended time interaction with the environment.
To do this, the agent must maximize return. For episodic tasks the return at time $t$ is defined as the sum of discounted rewards from $t$ to the end of the episode ($T$):
$G_{t}=r_{t}+\gamma\cdot r_{t+1}+\dots+\gamma^{T-t}r_{T}$
where $r_{t}$ is the reward observed at time $t$, and $\gamma$ is a discount between 0 and 1 on future rewards. This return definition can be written recursively:
$G_{t}$ $=r_{t}+\gamma\cdot G_{t+1}$
The action value function, which is used to choose actions, can be written as a function of the expected value of the return:
$Q(s,a)$ $=E\left[G_{t}\mid S_{t}=s,A_{t}=a\right]$
where $S_{t}$ is the state at time $t$ and $A_{t}$ is the action at time $t$.
There are two types of algorithm that can be used to learn how to behave in this type of environment: bootstrapping algorithms, which update the $Q$ value of each state and action based on the current estimate of the value of the next state(s) (e.g. Q-learning [14] and SARSA [7, 11] and non-bootstrapping algorithms, which work directly from samples of $G_{t}$ (e.g. Markov Chain Monte Carlo [2, 9]).
3 Background: Cooperative Multiagent RL
Multiagent Reinforcement learning often starts by considering an MDP that incorporates the states and actions of each agent in a set of agents. In the case of cooperating agents (the case closest to the problem we consider here), the agents are considered to be optimizing a joint reward function in this combined multi-agent MDP. Within this MDP, the number of states and actions grows exponentially with the number of agents. Work in multi-agent RL typically uses function approximation to simplify the large joint state space, or uses a simplified view of the joint state that is only partially observed by each agent or small group of agents, as in e.g. [5, 8].
In this work, we take a different approach. While in the future we will add function approximation, we begin by examining whether we need to consider the joint state and action space at all, in order to start with a problem formulation that is as simple as possible. We conclude that we do not need the joint state space, though we do need to consider the joint actions when making choices between them.
4 Basic Multiagent Algorithm Baseline
The problem we outline in this paper does not translate immediately to a standard tabular multiagent solution. In order to test against a realistic baseline, we have designed a multiagent solution with a basic communication protocol and a reward function that reflects the cooperative and competitive aspects of the task. As a multiagent problem, the task has the following properties: (1) Agents may only communicate when they split into multiple copies, in the start state. (2) Only the best agent to achieve the goal gets rewarded for doing so. (3) Costs are summed across all agents. This means that while the overall task is cooperative, individual agents compete to best achieve the goal and receive the reward for doing so.
To reflect these constraints, we implemented independent agents with communication and coordination in the start state. All agents receive negative cost rewards as they go, but only the best agent receives positive optimization reward at the end of the task.
Each agent remembers the joint action chosen at the start state. If $n$ agents choose the action set $\{a_{0},a_{1},\dots,a_{n}\}$ in the start state, then at every state $s$ in the episode we add this information to the state, forming the augmented state $(s,(a_{1},a_{2},\dots a_{n}))$. This provides enough information to estimate the agent’s probability of success in the race for the optimization reward in the context of the other agents, but increases the size of the state space. Once the state space is augmented in this way, each agent runs q-learning with a shared value function between all agents.
In the start state, when calculating the value of a candidate set of actions $(a_{1},a_{2},\dots a_{n})$, the values for each individual action in the context of the joint action are summed. The action combination with the highest value is chosen. From that point forward, the agents act independently, even if they are on similar paths.
5 Variance and Action Combinations
It may seem natural to order actions by their expected value when deciding which paths to send agents down. However, when combining multiple agents to find the highest performing one, actions with high variance may combine better than agents with low variance.
Consider a simple Multiarmed Bandit [12] problem with two arms, arm 1: normal distribution with $\mu=10,\sigma=1$ and arm 2: normal distribution with $\mu=5,\sigma=30$. Mean sample values for the maximum reward returned by combinations of up to 2 arms after 4000 sample, are highest when the noisiest arm is used twice, shown as follows:
Action 1 2 (1,1) (1,2) (2,2)
Est. Val. 10.0 5.22 10.56 19.70 22.11
Action (2, 2) is the best 2-action combination when the maximum of the two arms is used as the overall return. Due to it higher variance, it is more likely to generate combinations of outcomes that have at least one high return.
6 Multicopy Actions
We divide the reward function into two portions: cost, which reflects the additive costs of running more copies; and optimization, which reflects the improvements in success with additional agent copies. The algorithm should find the best trade-off between the two.
At certain states the agent may make up to $n$ copies of itself. In state $s$, considering multi-actions that include up to $n$ actions, the set of multi-action combinations available is:
[…]
where $\mathcal{A}(s)$ is the set of actions available at $s$. In some experiments we remove multi-actions with duplicate actions from consideration, while in others we include them.
The agent starts with a single copy, agent 00. If a copy $i$ chooses multiaction $m_{t}$ at time $t$, each agent copy produced is assigned a unique id. Define $ID_{t}$ to be the set of unique agent ids at time $t$, and $ID_{t}(i)$ to the the set of agent ids that are created from agent $i$ at time $t$. $m_{t}$ consists of a set of actions $a_{ij}$ indexed by the original agent $i$ and the agent $j$ which results from the action: $m_{t}=\{a_{ij}\mid j\in ID_{t}(i)\}$.
(a) Three Bridges Gridworld: quick, risky bridge (A); slow, safe bridge (B); medium bridge (C).
(b) Broken Bridges Gridworld with one broken bridge (A). The broken bridge is randomly chosen per episode.
Figure 1: Three-bridge gridworlds. Terminal states are marked with double squares. All other states have 4 actions: North, South, East, West. Postive terminal state rewards are optimization rewards, negative are costs. White squares have a per-step cost that is set for each experiment.
Return Definition for Multiple Actions. Our goal is to avoid using multi-sets of next states to tabulate the value function, and to use multi-sets of actions only in states where duplication of the agent is considered. We consider problems where the total sample return $G(t,i)$ for agent $i$ at time $t$ can be factored into two components, the cost return $G_{c}(t,i)$ and the optimization return $G_{o}(t,i)$:
$G(t,i)=G_{c}(t,i)+G_{o}(t,i)$ (1)
Costs are summed across all copies of the agent. Let $R_{c}(t,i)$ be the cost reward at time $t$ for agent $i$, and $G_{c}(t,i)$ be the cost return for agent $i$ at time $t$. The return for agent $i$ at time $t$ can be written in terms of the return for the set $ID_{t}(i)$ of agent copies created:
$G_{c}(t,i)$ $=\sum_{j\in ID_{t}(i)}\left[R_{c}(t,j)+\gamma G_{c}(t+1,j)\right].$ (2)
In episodic problems, different agent copies may finish at different times. $G_{c}$ for each agent ends at $T(i)$, the last timestep in which that agent is active:
$G_{c}(T(i),i)$ $=\sum_{j\in ID_{T(i)}(i)}R(T(i),j).$ (3)
Optimization rewards are only counted on the “best" path found. Therefore at each timestep they are maximized over the returns from the next timestep. Let $R_{o}(t,i)$ be the optimization reward at time $t$ for agent $i$, and $G_{o}(t,i)$ be the corresponding optimization return. When calculating overall optimization reward for the set of agent copies, we only count the best agent’s return:
$G_{o}(t,i)$ $=\max_{j\in ID_{t}(i)}\left[R_{o}(t,j)+\gamma G_{o}(t+1,j)\right].$ (4)
$G_{o}$ for agent $i$ ends at $T(i)$, the last timestep in which $i$ is active:
$G_{o}(T(i),i)$ $=\max_{j\in ID_{T(i)}(i)}R(T(i),j).$ (5)
Policy Evaluation for Multiple Actions. The value of multiaction $m$ in state $s$ is its expected value, at any time steps where $s,m$ is experienced:
$Q(s,$ $m)=E\left[G(t,i)\mid S_{t,i}=s,M_{t,i}=m\right]$
$=E\left[G_{c}(t,i)+G_{o}(t,i)\mid S_{t,i}=s,M_{t,i}=m\right]$
where $S_{t,i}$ is the state at time $t$ for agent $i$, and $M_{t,i}$ is the multi-action at time $t$ for agent $i$.
This expectation factors into the sum of the cost expectation and the optimization expectation:
$Q(s,m)=~{}$ $E\left[G_{c}(t,i)+G_{o}(t,i)\mid S_{t,i}=s,M_{t,i}=m\right]$
[…]
$E\left[G_{o}(t,i)\mid S_{t,i}=s,M_{t,i}=m\right]$
where $S_{t,i}$ is the state of agent $i$ at time $t$, and $M_{t,i}$ is its multiaction.
This factors into a value function for cost, $Q_{c}$, and a value function for optimization, $Q_{o}$:
$Q(s,m)$ $=Q_{c}(s,m)+Q_{o}(s,m)$
The cost q-function factors further, into a sum of q-values for individual actions $a\in m$:
$Q_{c}(s,m)$ $=E\left[G_{c}(t,i)\mid S_{t,i}=s,M_{t,i}=m\right]$
$=E\left[\sum_{j\in ID_{t}(i)}G_{c}(t,j)\mid S_{t,i}=s,\in M_{t,i}=m\right]$
If $A_{t,j}$ is defined as the individual action that produces agent copy $j$ at time $t$, then this becomes:
$=\sum_{a\in m}E\left[G_{c}(t,j)\mid S_{t,i}=s,A_{t,j}=a\right]$
$=\sum_{a\in m}Q_{c}(s,a)$
Therefore, for $Q_{c}$ we can use any learning technique, including bootstrapping methods like Q-learning [14, 12] or SARSA [7, 12] to learn the Q-function for individual actions, and sum for multiactions.
$Q_{o}$ cannot be simplified in the same way, as the expected value and maximization do not commute in general:
$Q_{o}(s,m)$ $=E\left[\max_{j\in ID_{t}(i)}G_{o}(t,j)\mid S_{t,i}=s,M_{t,j}=m\right]$
Here we must learn $Q_{o}$ for each combination of actions we consider. However, within the maximization step each $G_{o}$ may be calculated individually, allowing us to avoid a value function defined over the cross-product of states. If we are estimating $G_{o}$ over multiple times steps this does limit us to techniques which do not use bootstrapping like Markov Chain Monte Carlo [12].
Policy Improvement
When calculating the best action, the agent should optimize the total value:
$\pi$ $(s)=$
$\operatorname*{arg\,max}_{m}E\left[G_{c}(t,i)+G_{o}(t,i)\mid S_{t,i}=s,M_{t,i}=m\right].$
This function cannot be maximized separately over cost and optimization values, therefore action selection must be done using the joint value function and its policy.
7 Multicopy Algorithm
Multicopy agent
The multicopy agent combines a cost agent ($Q_{c}$), and an optimization agent ($Q_{o}$), and sums their values to calculate the joint multicopy policy $\pi$. This joint policy is used by both cost and optimization agents.
Cost and Optimization Agents
The cost agent estimates $Q_{c}$ for each state $s$ and individual action $a$ using Q-learning. To calculate the value of a multiaction $m$, it sums the values of its component actions. The optimization agent estimates $Q_{o}$ for each state $s$ and multiaction $m$, using Every Visit MCMC.
8 Gridworld Domains
The testing domains were chosen to examine the ability of the Multicopy agent to adapt the number of agent copies and the actions those copies take to improve performance under various levels of noise and cost.
Gridworld Environment. To model tasks with varying risk, delay and cost in a simple domain, we use a set of gridworld bridges of varying length and width, shown in Figure 1(a). This is an expanded version of a single bridge gridworld from the Pacman Projects [3]. Falling off a bridge represents task failure, and successfully making it across any bridge represents task success.
The three bridges present different trade-offs between speed and risk. Bridge (A) is short but narrow, useful when noise is low but risky when noise is high. Bridge (B) is long and wide: safer, but slower. Bridge (C) is a compromise.
Start state
In the start state (not shown), the agent may choose any single bridge or a combination of bridges, up to a maximum number of actions (set to 1, 2 or 3 per experiment). The actions in the start state have no noise, transitioning to the states labeled "A", "B" or "C" in Figure 1(a). An agent copy cannot transition back to the start state once it has chosen a bridge.
Noise
The gridworld includes an adjustable noise parameter [3]. With noise setting $\beta$, directional actions (N, S, E, W) on the bridges go in the intended direction with probability $1-\beta$, and go in the two orthogonal directions with probability $\beta/2$ each.
Optimization and Costs
The optimization reward, shown as positive reward in Figure 1(a), rewards successful crossings. The cost reward consists of a small negative reward on each step in the white bridge states, and a larger negative reward (shown) for falling off the bridge. The cost in bridge states varies per experiment.
(a) Basic learning Comparison: Noise 0.1
(b) Multiple Actions: Noise 0.1
(c) Long term convergence: Noise 0.1
(d) Basic learning Comparison: Noise 0.3
(e) Multiple Actions: Noise 0.3
(f) Long term convergence: Noise 0.3
Figure 2: Three Bridges Learning results for basic non-multicopy vs multicopy algorithms on the 3 bridges gridworld. Cost per step is -2 in all plots. Shaded lines are averaged over 50 trial runs, dark lines are rolling 30-episode averages of those values.
Broken Bridges
In the version of the three bridge gridworld shown in Figure 1(b), at the start of each episode one randomly chosen bridge is broken midway along (Bridge A in the diagram). All agents crossing the broken bridge during the episode will fail.
Exploration
On-policy techniques like MCMC can be overly cautious when noise is high. We therefore use a boltzmann distribution with action advantages and decreasing temperature over time, starting at 100, for exploration.
Algorithm Parameters
The learning rates for MCMC and Q-learning typically differ, with lower learning rates for MCMC. In all algorithms we use 0.05 for MCMC and 0.2 for Q-learning, linearly decreasing to 0 through the training episodes. We used a discount of 0.9 for all experiments.
9 Experiments
The experiments in this section aim to determine if the algorithm can find the right number of agent copies and combination of actions. First, we compare our algorithm from Section 7 to the basic algorithm with joint actions from Section 4 in the Three Bridges Gridworld, with various maximum action number settings. Then we compare the number and type of actions chosen for various cost and noise settings, to confirm that the algorithm is able to adjust the multiactions chosen according to the domain. Next we examine the difference in policy when we allow duplicate actions (more than one agent on the same bridge), and finally, some more nuanced experiments with duplicate actions.
Training and Testing
Each agent was trained for 7500 episodes and tested for 500 episodes, except where noted. During the testing phase, exploration and learning are turned off completely and only the optimal multiaction is chosen.
(a) Return. Per Episode return during testing phase for various noise and cost settings, after 7500 episodes of training. Final graph shows proportion improvement from the Joint Action return to the Multicopy return.
(b) Policy. Best actions in the start state for various noise and cost settings. Top action listed is chosen in most trials, shading shows proportion of trials in which this action is best. All actions best in $>20\%$ of trials are shown in decreasing order.
Figure 3: Varying Costs and Noise. Return and policy during testing for various cost and noise settings, showing the improvement in return from the Joint Action algorithm to Multicopy (a). Actions shown in (b) are best actions. 50 trial runs.
9.1 Learning Curves
Figure 2 contains learning curves for noise levels of 0.1 and 0.3, with bridge step-costs of -2. In Figure 2(a) and (d), we compare our multiagent baseline algorithm (Joint Actions) and our Multicopy algorithm with MCMC and Q-learning. For this comparison, the maximum number of actions is set to 1. Note that the joint action algorithm is similar to Q-learning (on-policy) while the multicopy algorithm is closer to MCMC (off-policy).
In Figure 2(b-c) and (e-f), we compare Multicopy and Joint Action algorithms with various maximum action settings. Figures (b) and (e) use the 7500 episode training time that we will use throughout the remainder of the paper, while Figures (c) and (f) use a longer training time of 50,000 episodes. The Multicopy algorithm does a fairly good job of learning as well in 7500 episodes as it does in 50,000, even with fairly high noise levels. The Joint Action algorithm, however, due to its fractured state space, does not learn as well when training time is short, particularly in (e), where noise is high.
(a) Return. Per Episode return during testing phase for various noise and cost settings.
(b) Policy. Best actions in the start state for various noise and cost settings. Top action listed is chosen in most trials, heatmap value is proportion of trials in which this action is best. All actions best in $>20\%$ of trials are shown in decreasing order.
Figure 4: Identical Duplicate Actions. Return and policy during testing for various cost and noise settings when identical duplicate actions are allowed. 50 trial runs.
9.2 Varying noise and costs
This section examines the Multicopy algorithm in a variety of cost/noise combinations, both to compare to the Joint Actions algorithm, and to examine what conditions lead to choosing different combinations of actions. In these initial experiments, duplicate actions are not allowed: only one agent may be sent on each of the three bridges. These experiments include limited training time of 7500 episodes.
As shown in Figure 3(a), the Multicopy algorithm generally performs better than the Joint Actions algorithm in these experiments. The highest proportional improvement is seen when noise is high (0.4) and costs are moderate (-4), and the lowest improvement when noise is 0.0.
Figure 3(b) details the policy chosen in the start state by each algorithm. The Joint Action algorithm is much more likely to choose a simple policy with 1 or 2 actions, perhaps because at this stage it has a better policy for getting across the bridges under those conditions.
The Multicopy algorithm has a more complex pattern of action choices. Generally one would expect that as cost magnitude increases, the agent should make do with fewer copies. Noise should have roughly the opposite effect: as noise increases, the agent should send more copies.
This rule of thumb largely applies for the upper and right rows of the graphs in Figure 3(b). When there is no noise, one action is consistently sufficient (first row). The area where three actions are best forms a wedge from the right side of the table, where costs are low, with a point at about -4, 0.2, where (A, B) and (A, B, C) have roughly the same value.
The number of actions goes back down to one when noise and costs are both high (the extreme lower left region). In this case the agent cannot reliably achieve the task even using all three bridges, and largely returns to using a single copy to keep costs down.
Between the regions where 3 actions are best, and the regions where 1 action is best, there are bands in which 2 actions is the optimal policy.
9.3 Duplicate Action Choices
In the initial experiments, the agent may only send a single agent on each of the three bridges, which limits the advantage to be gained from using all three agent copies. In this section we look at whether sending more than one agent on the same bridge can improve performance.
As in the previous section, we compare the performance of the Joint Action and Multicopy algorithms when a maximum of 3 actions are allowed in the start state.
Identical Actions
When we remove the requirement that there be only one agent per bridge in Figure 4, we see that the agent tends to choose to send multiple agents over the best bridge. In this case is also fairly noisy in its outcome due to its narrowness, so it also fulfills the criteria of Section 5. There is no correlation between the success or failure of the first agent and the second, so making multiple attempts at the same bridge can improve performance even if the agents do no coordination or cooperation on the bridge.
Multiple copies still does not help when there is no noise in the environment (the first rows of all heatmaps in Figure 4). While the Joint Action algorithm is closer in return to the Multicopy algorithm in these experiments (Figure 4(a)), it does make the odd choice to send three agent copies in this case. Generally the "best" action is not consistent across trial runs for this algorithm, however (see Figure 4(b)), indicating that it has not fully converged by this point.
For the Multicopy algorithm, Figure 4(b) shows fewer cases in which the agent chooses fewer than three copies of itself, though there are still a few cases where the agent chooses fewer copies to reduce costs.
9.4 Correlated Dangers
Thus far, we have examined a case in which the agent may send only one copy per bridge and a case in which the agent can and does send multiple copies on the same bridge. Next we examine a case in which despite having the option to send multiple copies on the same path, the rational behavior is to spread copies over different paths. One such scenario occurs when the noise or dangers on a path are correlated for all agents on the same path.
(a) Return. Per Episode return during testing phase for various noise and cost settings.
(b) Policy. Best actions in the start state for various noise and cost settings. Top action listed is chosen in most trials, heatmap value is proportion of trials in which this action is best. All actions best in $>20\%$ of trials are shown in decreasing order.
Figure 5: Broken Bridges Experiments. Return and policy during testing for various cost and noise settings when identical duplicate actions are allowed and a random bridge is broken in each trial. 20 trial runs.
We model this using the Broken Bridges gridworld in Figure 1(b). Here, each time the task is attempted, the agent finds that a random bridge is broken, and all agents on that same bridge fail at the task. This might correspond in more realistic domains to faults in parts of the wireless network, or the possibility of tunnel collapse in a multi-robot region for a robot navigating a collapsed building.
This appears to be a particularly difficult task for the Joint Actions agent to learn in the time allotted: in Figure 5(a) we can see that the Multicopy algorithm can do up to 12 times better in cases where noise is high. In Figure 5(b) we can see that this basic algorithm often still chooses only repetitions of bridge A, which is the shortest bridge and often the first that it learns to cross.
Since sending multiple agents on the same bridge can lead to both agents being lost, and the Multicopy agent, on the other hand, prefers to spread the copies over at least two different bridges to ameliorate risk, as shown in Figure 5(b).
10 Future Work
There are several research questions that arise from this work.
Wireless Networks
We will continue developing these algorithms in combination with modern function approximation methods such as Neural Networks [1, 15] in order to apply them to Mobile Wireless Networks [13, 10].
Other types of Optimization
In this paper, and in the Mobile Wireless Networks applications that inspired it, the agent seeks to optimize the amount of time required to complete a task. However, other types of optimization functions are possible within the same mathematical framework: building the best object in an assembly line, or getting the farthest in an exploration or rescue task.
Agent communication and coordination
In this work, we assume that communication between agents is expensive or unavailable much of the time, aside from the point at which the agent decides to duplicate itself. We therefore use a model where communication is limited to only the duplication step. This is essential in producing the simplified value function we use here. However, if possible, we will look for types of communication that are possible while maintaining this value function model.
Multiple goals
In some cases, an set of agents might need to achieve several similar goals at once (putting out several fires or finding several items). Part of our ongoing research includes adapting our algorithm to this case.
11 Conclusion
This paper makes several contributions. First, we have shown that for the specific multiagent problem with examine here, considering the joint state space of all agent copies is not necessary, and joint actions are necessary only when duplicating agents.
Second, we have shown that even using a simple very gridworld, the algorithm found a complex pattern of optimization that is easier to construct algorithmically than by hand-coding heuristics.
Third, we examined the cases in which the algorithm should use the same action more than once when creating agent copies, and cases in which it should use a diverse action set.
This formalism, while more restrictive than the general multiagent case, seems likely to be useful in Mobile Wireless Networks and other domains as well.